Please show me how to find the x intercept and y intercept of the following equation.
y=x^2-6x-7
Thanks so much!

Answer by stanbon(57203)

y=x^2-6x-7
To get the y-intercept, let x=0
Then y=0^2-6(0)-7=-7
Y-intercept is (0,-7)
To get the x-intercept, let y=0
Then x^2-6x-7=0
Factor to get:
(x-7)(x+1)=0
x=7 or x=-1
The X-intercepts are (7,0) and (-1,0)
Cheers,
stan H.